# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 3.2 A Sampler from Bounded Independence
$\ell\cdot\log\frac{1}{p}\cdot\widetilde{O}(d^{\prime})=\widetilde{O}\mathopen{}\mathclose{{}\left(\log^{2}\frac{1}{\delta_{\Gamma}}\cdot\frac{\log n}{\log m}}\right),$
and computing the sequences themselves takes $\ell\cdot\widetilde{O}(n)$ time. Observing that $\ell=O(\log m)$, the proof is concluded. ∎
We will need to somewhat extend Lemma 3.8 and use the simple, yet crucial, property of our bounded independence sampling: A subset of the coordinates of a $(b,\varepsilon)$-wise independent distribution with distinct samples is itself a $(b,\varepsilon)$-wise independent distribution with distinct samples. [^13] We note that the distinct-samples sampler given in [Vad04] does not seem to enjoy a similar property. Thus, if we wish to sample multiple times, say using $m_{1}\leq\ldots\leq m_{t}<n$ samples, we can use one sample from a sampler that outputs $m_{t}$ coordinates, and truncate accordingly to create the other samples. We only need to note that: (1) the sampling parameters are determined by the different $m_{i}$-s (and in particular, $m_{1}$ should be large enough), and (2) $m_{t}$ needs to be small enough compared to $n$, so that we can get enough distinct symbols. We summarize this observation in the next lemma.
@@
For any positive integers $n$ and $m_{1}<\ldots<m_{t}\leq n$, any $\delta\in(0,1)$ and any constant $\varepsilon$ such that $m_{t}\leq\frac{\varepsilon}{16}n$, there exists an explicit function $\Gamma\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\rightarrow[n]^{m_{t}}$ with $d=O\mathopen{}\mathclose{{}\left(\frac{\log n}{\log m_{1}}\cdot\log\frac{1}{\delta_{\Gamma}}}\right)$ that satisfies the following.
@@
    For any $i\in[t]$, the function $\Gamma_{i}$, that on input $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$ outputs $\Gamma(y)|_{[1,m_{i}]}$, is a $(\delta,\varepsilon)$ sampler, and each sample contains distinct symbols.
@@
    On input $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$, $\Gamma(y)$ can be computed in time $\widetilde{O}(n+\log^{2}(1/\delta))$.
@@ 3.3 Nearly-Linear Time Condensers
We first give the condenser based on multiplicity codes, due to Kalev and Ta-Shma [KT22].
Theorem 3.10 (the lossless KT condenser, [KT22]).
@@
For any constant $\alpha\in(0,1)$ the following holds, for every $n\in\mathbb{N}$, and any $0<\varepsilon\leq\frac{1}{n}$ and $k\geq\frac{256}{\alpha^{2}}\log^{2}\frac{n}{\varepsilon}$. There exists an explicit strong $(k,k^{\prime}=k+\ell,\varepsilon)$-condenser
@@
where $\ell=\mathopen{}\mathclose{{}\left(1+\frac{1}{\alpha}}\right)\log\frac{nk}{\varepsilon}=O_{\alpha}(\log\frac{1}{\varepsilon})$ and $m=(1+\alpha)k$. Note that the output entropy rate satisfies $\frac{k^{\prime}}{m}\geq 1-\alpha$. Moreover, given $x\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell}$, the output $\mathsf{KTCond}(x,y)$ can be computed in $\widetilde{O}(n)$ time.
In particular, if $\varepsilon^{\prime}=\sqrt{\varepsilon}$, then for all $(n,k)$-sources $X$ and a $(1-\varepsilon^{\prime})$-fraction of seeds $y$ it holds that $\mathsf{KTCond}(X,y)\approx_{\varepsilon^{\prime}}Z_{y}$, where $Z_{y}$ is an $(m=(1+\alpha)k,k^{\prime}-\ell=k\geq(1-\alpha)m)$-source. Note that the seed length is $\ell=O_{\alpha}(\log\frac{1}{\varepsilon^{\prime}})$.
@@
For the first part of the theorem statement we only need to establish the construction’s runtime. Given $x\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell}$, set a prime $p=\operatorname{poly}_{\alpha}(\frac{n}{\varepsilon})$, [^14] More precisely, they set $h=(2nk/\varepsilon)^{1/\alpha}$, and take $p$ to be a prime between $\frac{1}{2}h^{1+\alpha}$ and $h^{1+\alpha}$. and interpret $x$ as a polynomial $f_{x}\in\mathds{F}_{q}[X]$ of degree at most $d-1$, and $y$ as an element of $\mathds{F}_{p}$. Thus, $n=d\log q$, and we can safely ignore rounding issues, which can easily be addressed. The output $\mathsf{KTCond}(x,y)$ is the sequence of derivatives
$\mathopen{}\mathclose{{}\left(f(y),f^{\prime}(y),\ldots,f^{(m^{\prime})}(y)}\right),$
where $m^{\prime}=\frac{m}{\log q}$. By Lemma 2.1, computing the derivatives takes $\widetilde{O}(d)\cdot\log p=\widetilde{O}(n)$ time. The rest of the auxiliary operations are negligible compared to computing the derivatives.
@@
To see the “In particular” part of the theorem statement, fix an $(n,k)$-source $X$ and note that $Y\circ\mathsf{KTCond}(X,Y)\approx_{\varepsilon}Y\circ Z$ for some $Z$ such that $\mathbf{H}_{\infty}(Y\circ Z)\geq k^{\prime}$. Let $Z_{y}=(Z|Y=y)$. Then, an averaging argument gives that for a $(1-\sqrt{\varepsilon})$-fraction of seeds $y$ we have $\mathsf{RSCond}(X,y)\approx_{\sqrt{\varepsilon}}Z_{y}$. Since $Y$ is uniformly random over $\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell}$, we get that $\mathbf{H}_{\infty}(Z_{y})\geq k^{\prime}-\ell$, as desired. ∎

### v2

@@ 3.2 A Sampler from Bounded Independence
$\ell\cdot\log\frac{1}{p}\cdot\operatorname{polylog}(d^{\prime})=\operatorname{poly}\mathopen{}\mathclose{{\left(\log\frac{1}{\delta_{\Gamma}}\cdot\frac{\log n}{\log m}}}\right),$
and computing the sequences themselves takes $\ell\cdot\widetilde{O}(n)$ time, and observe that $\ell=O(\log m)$.
Finally, we need to argue that we can also handle the case where $n$ is not a prime power. One option to handle arbitrary $n$-s is to resort to almost $b$-wise independence. Specifically, we can start with a $\gamma$-biased distribution, with a small enough $\gamma$, over $n_{\mathsf{b}}=\lceil\tau^{-1}\log n\rceil n$ bits, and map each consecutive $\lceil\tau^{-1}\log n\rceil$ bits to $[n]$, for a small enough $\tau$, by simply taking the corresponding integer modulo $n$. We skip the details (to see the corresponding sampling property, see [XZ25]), and note that the result that uses the bounded-independence sampler, theorem˜4.1, only needs to invoke the sampler with $n$ being a prime power. [^18] In some more detail, in the proof of lemma 4.2 we use the $b$-wise distribution over $[n]$ in order to sub-sample from $X\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$. In the context of theorem 4.1, we can assume that $X$ has entropy rate $1-\alpha$ for an arbitrarily small constant $\alpha>0$, and the goal is to retain its high entropy rate when sub-sampling. Letting $p$ be the largest prime smaller than or equal to $n$ (which satisfies $p\geq n/2$ by Bertrand’s postulate), we have that $X^{\prime}=X_{[1,p]}$ has entropy rate at least $1-2\alpha$, and one can verify that sampling from $X^{\prime}$ works equally well, up to negligible loss in parameters. Moreover, we can find $p$ in time $\widetilde{O}(n)$ using, e.g., the deterministic AKS primality test [AKS04] that runs in time $\operatorname{polylog}n$ applied to each integer in $[n/2,n]$.
We will need to somewhat extend lemma˜3.8 and use the simple, yet crucial, property of our bounded independence sampling: A subset of the coordinates of a $b$-wise independent distribution with distinct samples is itself a $b$-wise independent distribution with distinct samples. [^19] We note that the distinct-samples sampler given in [VAD04], an instantiation of which we use in lemma 2.20, does not seem to enjoy a similar property. The advantage of lemma 2.20 over lemma 3.9 that will appear soon, is that it has better seed length. Thus, if we wish to sample multiple times, say using $m_{1}\leq\ldots\leq m_{t}<n$ samples, we can use one sample from a sampler that outputs $m_{t}$ coordinates, and truncate accordingly to create the other samples.
@@
For any positive integers $n$ and $m_{1}<\ldots<m_{t}\leq n$, any $\delta\in(0,1)$ and any constant $\varepsilon$ such that $m_{t}\leq\frac{\varepsilon}{16}n$, there exists an explicit function $\Gamma\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\rightarrow[n]^{m_{t}}$ with $d=O\mathopen{}\mathclose{{\left(\frac{\log n}{\log m_{1}}\cdot\log\frac{1}{\delta}}}\right)$ that satisfies the following.
@@
    For any $i\in[t]$, the function $\Gamma_{i}$, that on input $y\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$ outputs $\Gamma(y)|_{[1,m_{i}]}$, is a $(\delta,\varepsilon)$-averaging sampler, and each sample contains distinct symbols.
@@
    On input $y\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$, $\Gamma(y)$ can be computed in time $\widetilde{O}(n)+\operatorname{poly}(\log(1/\delta),(\log n)/(\log m_{1}))$.
Proof Sketch.
Let $\Gamma\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\rightarrow[n]^{m_{t}}$ be the $(\delta,\varepsilon)$-averaging sampler of lemma˜3.8, set with $\delta_{\Gamma}=\delta$ and $\eta=\varepsilon/2$. Fix some $i\in[t]$, and let $\Gamma_{i}$ be as described in the lemma’s statement. Inspecting the proof of lemma˜3.8, and using the same notation, we see that the output of $\Gamma_{i}$ can be obtained by truncating the output of $\mathsf{BI}_{b}$ to the first $m^{\prime}_{i}=(1+\eta)m_{i}$ bits and using the same expander random walk in order to get $m_{i}$ distinct symbols with high probability. Note that:
  1. 1.
    (sampling) $X_{[1,m^{\prime}_{i}]}$ is a $b$-wise independent distribution, whose sampling properties are determined by $m^{\prime}_{i}$. We then need to make sure that the smallest $m^{\prime}_{i}$ is large enough, by setting $b=O\mathopen{}\mathclose{{\left(\frac{\log(1/\delta_{\Gamma})}{\log m_{1}}}}\right)$, as in the proof of lemma˜3.8.
  2. 2.
    (distinctness) In order for the probability that $X_{[1,m^{\prime}_{i}]}$ has at least $m_{i}$ distinct symbols to be large enough (recall that $X$ itself is over $m^{\prime}_{t}$ symbols), $m_{i}\leq m_{t}$ needs to be small enough compared to $n$. And indeed, we take $m_{i}\leq\frac{\eta}{8}n=\frac{\varepsilon}{16}n$.
Once the two properties above are guaranteed, we can amplify the probability to sample a string with sufficiently many distinct symbols via expander random walks, exactly as in the proof of lemma˜3.8. ∎
@@ 3.3 Nearly-Linear Time Condensers
In this section we argue that modern constructions of condensers with nearly-optimal parameters [GUV09, KT22] are computable in nearly-linear time, possibly after a reasonable one-time preprocessing step. We will repeatedly use the fact that these condensers allow us to transform, in time $\widetilde{O}(n)$ and using nearly-optimal seed length $O(\log(n/\varepsilon))$, an arbitrary input $(n,k)$-source $X$ into an output $\varepsilon$-close to a source $X^{\prime}$ of length $m\approx k$ and min-entropy rate $1-\alpha$, for any constant $\alpha\in(0,1)$ of our choice. We provide formal statements below, without being explicit about the precise relationship between the parameter $\alpha$ and the various constants in the seed length and entropy requirement for the sake of readability. More precise control over the constants in these condensers can be found, for example, in [KT22, Theorems 1 and 2].
3.3.1 The Lossless KT Condenser
We first give the lossless KT condenser based on multiplicity codes over $\mathds{F}_{q}$, due to Kalev and Ta-Shma [KT22]. This condenser works when the min-entropy requirement $k=\Omega(\log^{2}(n/\varepsilon))$. Then, we discuss how it can be converted to a condenser over bits with only a negligible loss in parameters.
Theorem 3.10 (the lossless KT condenser over $\mathds{F}_{q}$ [KT22, adapted]).
@@ 3.3.1 The Lossless KT Condenser
For any constant $\alpha\in(0,1)$ there exist constants $C_{\alpha},C^{\prime}_{\alpha}>0$ such that the following holds for every $n\in\mathbb{N}$, $\varepsilon>0$, $k\geq C_{\alpha}\log^{2}(n/\varepsilon)$, and a prime power $q=p^{r}$ with $p\geq n$ satisfying $\frac{C^{\prime}_{\alpha}}{2}\log(n/\varepsilon)\leq\log q\leq C^{\prime}_{\alpha}\log(n/\varepsilon)$. There exists a strong $(k,k^{\prime}=k+\log q,\varepsilon)$-condenser
$\mathsf{KTCond}^{\prime}\colon\mathds{F}_{q}^{n^{\prime}}\times\mathds{F}_{q}\rightarrow\mathds{F}_{q}^{m^{\prime}}$
where $n^{\prime}=\mathopen{}\mathclose{{\left\lceil\frac{n}{\lfloor\log q\rfloor}}}\right\rceil$ and $m^{\prime}\leq(1+\alpha)\frac{k}{\log q}$. Moreover, given $p$, an irreducible polynomial over $\mathds{F}_{p}$ of degree $r$, $x\in\mathds{F}_{q}^{n}$, and $y\in\mathds{F}_{q}$, the output $\mathsf{KTCond}^{\prime}(x,y)$ can be computed in $\widetilde{O}(n^{\prime}\log q)=\widetilde{O}(n)$ time.
Proof.
We only need to establish the construction’s runtime. Given $x\in\mathds{F}_{q}^{n^{\prime}}$ and $y\in\mathds{F}_{q}$, interpret $x$ as a polynomial $f\in\mathds{F}_{q}[X]$ of degree at most $n^{\prime}-1$. The output $\mathsf{KTCond}(x,y)$ is the sequence of derivatives
$\mathopen{}\mathclose{{\left(f(y),f^{\prime}(y),\ldots,f^{(m^{\prime}-1)}(y)}}\right).$
By lemma˜2.1, computing the derivatives takes $\widetilde{O}(n^{\prime})\cdot\log q=\widetilde{O}(n^{\prime}\log q)=\widetilde{O}(n)$ time since we are also given $p$ and an irreducible polynomial over $\mathds{F}_{p}$ of degree $r$. The rest of the auxiliary operations are negligible compared to computing the derivatives. ∎
Remark 3.11 (the KT condenser preprocessing step).
The KT condenser must be instantiated with an appropriate field size $q=\operatorname{poly}(n/\varepsilon)$ and requires performing operations over $\mathds{F}_{q}$. Hence, as stated in theorem˜3.10, we need to know the characteristic $p$ of $\mathds{F}_{q}$ and an irreducible polynomial over $\mathds{F}_{p}$ of the appropriate degree. Since these objects only need to be generated once for a given set of extractor parameters $(n,k,\varepsilon,m)$, we view this is as a one-time preprocessing step. Our choice of $q$ influences the complexity of constructing a model of $\mathds{F}_{q}$. For example, if we aim for prime field size, then we need to generate a prime $q=\operatorname{poly}(n/\varepsilon)$. We know how to do that in randomized time $\operatorname{polylog}(n/\varepsilon)$ (e.g., see [SHO05, Section 9.4]). We believe that this is already reasonable when seen as a one-time preprocessing step, because generating large primes is a well-studied problem and practical randomized algorithms exist.
Nevertheless, we can do even better and avoid the preprocessing step if $\varepsilon$ is not very small by exploiting the fact that the KT condenser works with any prime power $q=p^{r}=\operatorname{poly}(n/\varepsilon)$ for any prime $p\geq n$ [KT22, Theorem 3]. First, a prime $p\in[n,2n]$ is guaranteed to exist by Bertrand’s postulate, and we can find it deterministically in time $\widetilde{O}(n)$. [^20] The procedure that sequentially tests the primality of all integers in $[n,2n]$ is already sufficiently quick. Each primality test takes time $\operatorname{polylog}(n)$ using, for example, the AKS primality test [AKS04], and so the overall complexity is $\widetilde{O}(n)$. Second, a deterministic algorithm of Shoup [SHO90, Theorem 3.2] finds an irreducible polynomial of degree $r$ over $\mathds{F}_{p}$ in time $\widetilde{O}(\sqrt{p}\cdot r^{4+\delta})$ for any constant $\delta>0$. Since $p\leq 2n$ and $r=\log_{p}(\operatorname{poly}(n/\varepsilon))=O(\log(n/\varepsilon))$, we conclude that Shoup’s algorithm runs in time $\widetilde{O}(n)$ provided that $\varepsilon\geq 2^{-Cn^{\gamma}}$ for any constant $\gamma<1/8$ and a sufficiently large constant $C>0$. In particular, it suffices to have, say, $\varepsilon\geq 2^{-Cn^{0.1}}$.
In sum, the preprocessing step for the KT condenser requires either (1) randomized $\operatorname{polylog}(n/\varepsilon)$ time, or (2) deterministic $\widetilde{O}(n+\sqrt{n}\cdot\log^{4+\delta}(1/\varepsilon))$ time. In particular, this leads to the distinction between items˜1 and 2 in ˜1.
The condenser from theorem˜3.10 receives and outputs vectors over $\mathds{F}_{q}$. We would like to have a version of this condenser that works with vectors over bits. Towards that end, for completeness, we formally state and prove a (standard) transformation.
Lemma 3.12.
Suppose that $\mathsf{Cond}_{0}\colon\mathds{F}_{q}^{n^{\prime}}\times\mathds{F}_{q}\to\mathds{F}_{q}^{m^{\prime}}$ is a strong $(k,k^{\prime},\varepsilon)$-condenser. Then, there exists $\mathsf{Cond}\colon\{0,1\}^{n}\times\{0,1\}^{\ell}\to\{0,1\}^{m}$ with $n=n^{\prime}\cdot\lfloor\log q\rfloor$, $\ell=\lfloor\log q\rfloor$, and $m=m^{\prime}\cdot\lceil\log q\rceil$ that is a strong $(k+\ell,k^{\prime},3\sqrt{\varepsilon})$-condenser. Furthermore, if $\mathsf{Cond}_{0}$ is computable in time $T$ then $\mathsf{Cond}$ is computable in time $T+\widetilde{O}(n)$.
Proof.
The condenser $\mathsf{Cond}$ works as follows. Given an input $x\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$, we interpret it as an element of $\mathds{F}_{q}^{n^{\prime}}$ by mapping each consecutive block of $\lfloor\log q\rfloor$ bits to an $\mathds{F}_{q}$-element and discarding up to $\lceil\frac{n}{\lfloor\log q\rfloor}\rceil\cdot\lfloor\log q\rfloor-n\leq\lfloor\log q\rfloor$ bits at the end of $x$. Denote this mapping by $\psi\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\to\mathds{F}_{q}^{n^{\prime}}$. Given a seed $y\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell}$, we interpret it as an $\mathds{F}_{q}$-element, since $\ell=\lfloor\log q\rfloor$. Denoting this mapping by $\phi\colon\mathds{F}_{2}^{\ell}\rightarrow\mathds{F}_{q}$, we have that $\mathbf{H}_{\infty}(\phi(U_{\ell}))\geq\ell$.
Equipped with an $\mathds{F}_{q}^{n^{\prime}}$-element $x$, and an $\mathds{F}_{q}$-element $y$, we are ready to compute $\mathsf{Cond}_{0}(x,y)\in\mathds{F}_{q}^{m^{\prime}}$. Mapping the output into bits is done similarly, by writing the binary encoding of each $\mathds{F}_{q}$-element using $\lceil\log q\rceil$ bits. There are a few possible losses along the way:
  1. 1.
    We get an $(n,k)$ source $X$, but feed the condenser a source $X^{\prime}\sim\mathds{F}_{q}^{n^{\prime}}$ with entropy at least $k-\lfloor\log q\rfloor=k$-$\ell$.
  2. 2.
    We do not use a uniform seed, but rather a seed that lacks at most $1$ bit of entropy.
  3. 3.
    Mapping $\mathsf{Cond}_{0}(\psi(x),\phi(y))$ to the binary $\mathsf{Cond}(x,y)$ may increase the output by $m^{\prime}$ non-entropic bits. This simply increases the output length slightly, but not the output entropy.
To handle (1), we simply increase the entropy of the source. To handle (2), we prove that strong condensers work with entropy-deficient random seeds, as long as the error is good enough.
Claim 3.13.
Let $\mathsf{Cond}\colon\mathds{F}_{q}^{n}\times\mathds{F}_{q}\rightarrow\mathds{F}_{q}^{m}$ be a strong $(k,k^{\prime},\varepsilon)$ condenser, and let $Y$ be a random variable with min-entropy $\log(q)-g$. Then, $(Y,\mathsf{Cond}(X,Y))$ is $\varepsilon^{\prime}$-close to a random variable with min-entropy $k^{\prime}$, where $\varepsilon^{\prime}=(2^{g}+1)\sqrt{\varepsilon}$.
Proof.
Let $X$ be a random variable over $\mathds{F}_{q}^{n}$ with min-entropy $k$. By an averaging argument, we know that there exists a set $\mathbf{B}\subseteq\mathds{F}_{q}$ of size $|\mathbf{B}|\leq\sqrt{\varepsilon}\cdot q$ such that for any $y\notin\mathbf{B}$ it holds that $\mathsf{Cond}(X,y)$ is $\sqrt{\varepsilon}$-close to a random variable with min-entropy $k^{\prime}-\log q$. But now,
$\Pr[Y^{\prime}\in\mathbf{B}]=\sum_{y\in\mathbf{B}}\Pr[Y=y]\leq|\mathbf{B}|\cdot 2^{-(\log(q)-g)}\leq\sqrt{\varepsilon}\cdot 2^{g}.\qed$
In our case we invoke ˜3.13 with $g=1$, so we can simply set the new error parameter $\varepsilon^{\prime}$ to be $\varepsilon^{\prime}=\varepsilon^{2}/3$, where $\varepsilon$ is the designated distance to high min-entropy. A bit more formally, guaranteeing that $(U_{d},\mathsf{Cond}(X,U_{d}))$ is $\varepsilon^{\prime}$-close to entropy $k^{\prime}$ implies that $(\phi(U_{d}),\mathsf{Cond}(X,\phi(U_{d})))$ is $\varepsilon$-close to entropy $k^{\prime}$ as well. Finally, for the running time claim note that computing the mappings $\psi$ and $\phi$ and converting the output in $\mathds{F}_{q}^{m^{\prime}}$ to $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$ can be done in time $\widetilde{O}(n)$. ∎
Combining theorem˜3.10 and lemma˜3.12 yields the following.
Theorem 3.14 (the KT condenser over bits [KT22]).
For any constant $\alpha\in(0,1)$ there exist constants $C_{\alpha},C^{\prime}_{\alpha}>0$ such that the following holds for every $n\in\mathbb{N}$, $\varepsilon>0$, and $k\geq C_{\alpha}\log^{2}(n/\varepsilon)$. There exists a strong $(k,k^{\prime},\varepsilon)$-condenser
@@
where $\ell\leq C^{\prime}_{\alpha}\log(n/\varepsilon)$, $m\leq(1+\alpha)k$, and $k^{\prime}=k\geq\ell+(1-\alpha)m$. Moreover,
  • Given $x\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell}$, a prime power $q=p^{r}$ in $2^{\ell},2n\cdot 2^{\ell}]$ with $p\in[n,2n]$ and an irreducible polynomial over $\mathds{F}_{p}$ of degree $r$, the output $\mathsf{KTCond}(x,y)$ can be computed in $\widetilde{O}(n)$ time.
  • Given only $x\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell}$, $\mathsf{KTCond}(x,y)$ can be computed in $\widetilde{O}(n+\sqrt{n}\cdot\log(1/\varepsilon)^{5})$ time. Note that this is $\widetilde{O}(n)$ when $\varepsilon\geq 2^{-Cn^{0.1}}$.
In particular, if $\varepsilon^{\prime}=\sqrt{\varepsilon}$ and $C_{\alpha}$ is chosen large enough compared to $C^{\prime}_{\alpha}$, then for all $(n,k)$-sources $X$ and a $(1-\varepsilon^{\prime})$-fraction of seeds $y$ it holds that $\mathsf{KTCond}(X,y)\approx_{\varepsilon^{\prime}}Z_{y}$, where $Z_{y}$ is an $(m,k^{\prime}-\ell\geq(1-\alpha)m)$-source.
@@
Fix a target $\alpha\in(0,1)$, and let $\beta\in(0,1)$ be a constant to be set appropriately small depending on $\alpha$. We invoke theorem˜3.10 with $\beta$ in place of $\alpha$ and $q$ a prime such that $\frac{C^{\prime}_{\beta}}{2}\log(n/\varepsilon)\leq\log q\leq C^{\prime}_{\beta}\log(n/\varepsilon)$ to get $\mathsf{KTCond}^{\prime}\colon\mathds{F}_{q}^{n^{\prime}}\times\mathds{F}_{q}\to\mathds{F}_{q}^{m^{\prime}}$, a strong $(k-\log q,k^{\prime}=k,\varepsilon^{2}/3)$-condenser with $n^{\prime}=\mathopen{}\mathclose{{\left\lceil\frac{n}{\lfloor\log q\rfloor}}}\right\rceil$ and $m^{\prime}\leq(1+\beta)\frac{k}{\log q}$. Then, we apply lemma˜3.12 to $\mathsf{KTCond}^{\prime}$. This yields our $\mathsf{KTCond}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell}\to\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$, a strong $(k,k^{\prime}=k,\varepsilon)$-condenser with $\ell=\lfloor\log q\rfloor=\Theta_{\beta}(\log(n/\varepsilon))$ and
$m=m^{\prime}\cdot\lceil\log q\rceil\leq(1+\beta)k+m^{\prime}\leq(1+\beta)k+(1+\beta)k/\ell\leq(1+2\beta)k\leq(1+\alpha)k,$
provided that $\beta\leq\alpha/2$. Furthermore, if $C_{\beta}$ is set sufficiently larger than $C^{\prime}_{\beta}$, then $k^{\prime}-\ell=k-\ell\geq(1-\beta)k$, and so
$\frac{k^{\prime}-\ell}{m}\geq\frac{(1-\beta)k}{(1+2\beta)k}\geq 1-\alpha,$
provided that $\beta>0$ is sufficiently smaller than $\alpha$. The first part of the theorem statement follows if we set the new $C_{\alpha}$ and $C^{\prime}_{\alpha}$ to be $C_{\beta}$ and $C^{\prime}_{\beta}$, respectively. For the running time, note that $\mathsf{KTCond}^{\prime}$ is computable in time $\widetilde{O}(n)$ given $p$ and an irreducible polynomial over $\mathds{F}_{p}$ of degree $r$, and the transformation in lemma˜3.12 still runs in time $\widetilde{O}(n)$. The remaining claim about the running time given only $x$ and $y$ was discussed in remark˜3.11.
@@
To see the “In particular” part of the theorem statement, fix an $(n,k)$-source $X$ and note that $(Y,\mathsf{KTCond}(X,Y))\approx_{\varepsilon}(Y,Z)$ for some $Z$ such that $\mathbf{H}_{\infty}(Y,Z)\geq k^{\prime}$. Let $Z_{y}=(Z|Y=y)$. Then, an averaging argument gives that for a $(1-\sqrt{\varepsilon})$-fraction of seeds $y$ we have $\mathsf{KTCond}(X,y)\approx_{\sqrt{\varepsilon}}Z_{y}$. Since $Y$ is uniformly random over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell}$, we get that $\mathbf{H}_{\infty}(Z_{y})\geq k^{\prime}$-$\ell$, as desired. ∎
3.3.2 The Lossy RS Condenser